OpenStudy (anonymous):

Use the function rule f(x)=5x+1. (Find each output) f(3)

OpenStudy (anonymous):

5(3)+1=16 f(x)=16

OpenStudy (ash2326):

We have \[f(x)=5x+1\] We have to find f(3), so put x=3 in f(x) \[f(3)=5\times 3+1=16\]